A consultant traveled 5 hours to attend a meeting. The return trip took only 4 hours because the speed was 5 miles per hour faster. What was the consultant's speed each way?
Answer by funmath(2933)

A consultant traveled 5 hours to attend a meeting. The return trip took only 4 hours because the speed was 5 miles per hour faster. What was the consultant's speed each way?
Distance = rate * time (d=rt)
The distance is the same both ways so:
rate to mtg* time to mtg= rate from mtg*time from mtg
Let the rate to mtg be: r
The time is:5
The distance to mtg is: 5r
Then the rate from mtg is: r+5
The time from mtg is: 4
The distance from mtg is: 4(r+5)=4r+20
Problem to solve:
5r=4r+20
5r-4r=4r-4r+20
r=20
The rate to mtg is: r=20 mph
The rate from mtg is: r+5=20+5=25 mph
Happy Calculating!!!!